# Instance-Adaptive Keypoint Learning with Local-to-Global Geometric Aggregation for Category-Level Object Pose Estimation

### v2

@@ IV-B2 Results on CAMERA25
Table II reports the performance of INKL-Pose compared with state-of-the-art methods on the CAMERA25 dataset. As shown in the table, INKL-Pose achieves outstanding accuracy across nearly all evaluation metrics. Although its IoU50 metric reaches 94.4%, slightly lower than SpherePose [29] (achieving 94.8%), it significantly outperforms all existing approaches on the remaining metrics, especially under the more rigorous pose accuracy thresholds. Specifically, compared to prior-based methods, INKL-Pose maintains a clear advantage without relying on pre-defined shape priors. In addition, when compared with AG-Pose [14], the current state-of-the-art shape-prior-free method, INKL-Pose achieves notable gains of 1.5% and 1.3% under the strict 5°2cm and 5°5cm metrics, respectively. These results, along with the strong performance observed on the REAL275 dataset, provide compelling evidence of the robustness and effectiveness of our proposed local-to-global geometric feature aggregation framework.
IV-B3 Results on HouseCat6D
To further validate the generalization capability of the proposed INKL-Pose, we conduct experiments on the more challenging real-world dataset HouseCat6D [40]. As reported in Table III, INKL-Pose achieves state-of-the-art performance across all evaluation metrics. Compared with the current state-of-the-art AG-Pose [14], our method surpasses it under all evaluation metric, with gains of 3.9% on 5°2cm, 5.3% on 5°5cm, 6.2% on 10°2cm, and 7.6% on 10°5cm. Moreover, Fig. 5 presents a detailed per-category comparison of 3D IoU, rotation, and translation accuracy between INKL-Pose and AG-Pose. The results reveal that INKL-Pose outperforms AG-Pose on the majority of metrics, particularly excelling in rotation estimation. Furthermore, Fig. 6 presents qualitative results of the predicted poses, further showcasing the strong generalization ability of INKL-Pose in complex real-world scenarios.
Additionally, Fig. 2 offers intuitive visual comparisons between INKL-Pose and other state-of-the-art methods on both the REAL275 and HouseCat6D datasets. These results collectively demonstrate the superior generalization capability and practical applicability of our method in real-world scenarios.
@@ IV-B3 Results on HouseCat6D
Figure 5: Per-category comparison between INKL-Pose and AG-Pose[14] on the HouseCat6D dataset[40], including mean Average Precision (mAP) for 3D IoU, rotation, and translation.
Figure 6: Visualization results of INKL-Pose on the HouseCat6D dataset [40]. Green bounding boxes denote ground truth, while red bounding boxes represent the predictions from our INKL-Pose. The predicted boxes show high consistency with the ground truth.
@@ IV-C Ablation Studies
In this section, we conduct comprehensive ablation studies to evaluate the contributions of key components in the proposed INKL-Pose. All experiments are performed on the REAL275 [10] dataset.
Figure 7: Performance comparison between INKL-Pose and state-of-the-art methods on the REAL275 and HouseCat6D real-world datasets [^2] This figure was inspired by the drawing style of CleanPose [30]. We gratefully acknowledge the authors for their work..
@@ IV-C1 Effect of the Number of Keypoints
IV-C2 Effect of Keypoint Supervision
In the Instance-Adaptive Keypoint Generator (IAKG) module (Section III-C), we introduce two auxiliary losses: $\mathcal{L}_{surf}$ and $\mathcal{L}_{sep}$, to promote uniform surface coverage and enforce spatial diversity among keypoints. To validate the effectiveness of this loss design, we conduct two types of experiments: (1) an ablation study by selectively removing each loss term, and (2) a comparative study by replacing our losses with those used in AG-Pose [14]. As shown in Table IV, removing $\mathcal{L}_{surf}$ results in a 2.8% drop under the 5°2cm metric, demonstrating its role in promoting even keypoint distribution and suppressing outliers. In contrast, removing $\mathcal{L}_{sep}$ leads to a larger performance degradation of 16.3%, highlighting its importance in preventing keypoint clustering. When both losses are removed, the performance drops sharply by 44.7%, confirming the necessity of the proposed supervision strategy. In the comparative setting, we replace our loss functions with those adopted in AG-Pose, which include an object-centric distribution loss and a global diversity term. Compared to these alternatives, our proposed losses provide stronger and more stable geometric supervision. Specifically, $\mathcal{L}_{surf}$ leverages bidirectional Chamfer Distance to ensure comprehensive surface coverage, while AG-Pose employs a one-sided nearest-distance constraint. Meanwhile, $\mathcal{L}_{sep}$ introduces localized repulsion based on nearest neighbors, enabling finer-grained control over spatial dispersion than the global pairwise formulation used in AG-Pose. Together, these results demonstrate the superiority of our loss design in guiding keypoint learning with both semantic consistency and geometric structure.
IV-C3 Effect of Positional Embeddings
We further analyze the contribution of positional information in enhancing feature representation. In INKL-Pose, positional encodings are introduced at two distinct stages: the embedding $\mathcal{E}_{\mathcal{P}}$ in the Pointwise Feature Encoder (Section III-B), which compensates for lost absolute position cues, and $\mathcal{E}_{kpt}$ in the LKFA module (Section III-D), which improves local spatial awareness. As reported in Table IV, removing either embedding leads to a significant performance drop, highlighting their complementary roles in facilitating accurate geometric perception.
IV-C4 Effect of the LKFA and GKFA Modules
We evaluate the individual and combined contributions of the LKFA and GKFA modules introduced in Section III-D and Section III-E, respectively. As shown in Table IV, removing the LKFA causes a 2.1% drop in the 5°2cm metric, primarily due to the loss of fine-grained geometric features. In contrast, removing the GKFA results in a larger decline of 3.0%, indicating its role in maintaining global consistency. When both modules are removed, performance drops by 3.7%, demonstrating their complementary importance in accurate pose estimation.
IV-C5 Effect of the Number of LKFA and GKFA Modules
We investigate the impact of the number of stacked keypoint feature aggregation blocks, where each block comprises both a Local Keypoint Feature Aggregator (LKFA) and a Global Keypoint Feature Aggregator (GKFA). As shown in Table IV, increasing the number of blocks from $S=4$ to $S=12$ consistently improves performance across all evaluation metrics. These results demonstrate that deeper cascaded aggregation enhances both local and global keypoint representations by progressively refining fine-grained geometric details and capturing long-range structural dependencies. However, further increasing the number of blocks to $S=16$ results in performance degradation, as excessive stacking introduces feature redundancy and oversmoothing, which reduce keypoint distinctiveness and impair pose estimation accuracy. Overall, $S=12$ offers an optimal balance between accuracy and efficiency, effectively capturing both local geometry and global structure without incurring unnecessary complexity.
IV-C6 Effect of Neighborhood Size
To evaluate the influence of neighborhood size $K$ in the LKFA module, we vary the number of nearest neighbors used during local feature aggregation. As shown in Table IV, when $K=1$, the model only uses the center point feature, which limits its ability to capture surrounding geometric and semantic context. Increasing $K$ enhances local structure modeling, with optimal performance achieved at $K=4$. Further increasing $K$ introduces redundant information, leading to no additional gains.
IV-C7 Effect of Bidirectional Mamba and Feature Sequence Flipping
We also examine the contributions of the bidirectional Mamba architecture and the proposed Feature Sequence Flipping (FSF) strategy within the GKFA module (Section III-E), with results summarized in Table IV. As the results show, replacing the bidirectional Mamba with its unidirectional counterpart (Uni-Mamba) leads to consistent performance degradation, confirming the effectiveness of bidirectional modeling for global feature aggregation. Furthermore, we compare FSF with the commonly used Point Sequence Flipping (PSF) strategy. PSF amplifies the non-causal nature of point clouds by reversing the sequence along spatial indices, thereby disrupting the alignment between feature order and geometric structure. This misalignment leads to a clear performance drop, even falling below the unidirectional baseline. In contrast, FSF retains the original point order and achieves bidirectional information flow in the feature space, significantly improving robustness to the unordered and non-causal nature of point cloud data.
TABLE IV: Ablation studies on the REAL275 dataset[10]. Default configurations are colored in gray, with best results highlighted in bold.
@@ IV-C7 Effect of Bidirectional Mamba and Feature Sequence Flipping
(a) Effect of the number of keypoints.
@@
Full ($\mathcal{L}_{sep}$+$\mathcal{L}_{surf}$) 63.2 68.3 80.9 88.0
w/ $\mathcal{L}_{sep}$ 60.8 67.7 77.8 86.2
w/ $\mathcal{L}_{surf}$ 46.9 53.0 69.1 78.8
w/ AG-Pose losses 60.9 67.0 79.9 87.2
@@
(b) Effect of keypoint supervision.
@@
(c) Effect of positional embeddings.
@@
None 59.5 67.7 78.0 88.3
(d) Effect of the LKFA and GKFA modules.
@@
(e) Effect of the number of LKFA and GKFA modules.
@@
(f) Effect of neighborhood size.
@@
(g) Effect of bidirectional Mamba and feature sequence flipping.
@@ V Conclusion
In this paper, we propose INKL-Pose, a novel keypoint-based framework for category-level object pose estimation that enhances generalization to unseen object instances through instance-adaptive keypoint learning and local-to-global geometric aggregation. Specifically, we introduce an Instance-Adaptive Keypoint Generator that predicts semantically consistent and geometrically informative keypoints, along with two complementary modules for feature refinement: a Local Keypoint Feature Aggregator that captures fine-grained geometric details and a Global Keypoint Feature Aggregator based on bidirectional Mamba with a Feature Sequence Flipping strategy for structure-aware global perception. Furthermore, we design a surface loss and a separation loss to explicitly supervise keypoint distribution, providing strong geometric guidance for robust pose estimation. Extensive experiments on CAMERA25, REAL275, and HouseCat6D demonstrate that INKL-Pose achieves state-of-the-art performance across both synthetic and real-world datasets, significantly outperforming existing approaches. In future work, we will focus on extending our framework to dynamic and temporal settings, enabling sequential pose tracking, and explore uncertainty-aware keypoint modeling to further improve robustness under occlusion and sensor noise.
Acknowledgments
This work was supported by the Wuhan Young Talent Program (Wuhan Talent Initiative) (Grant No. X24D090), the Hubei Science and Technology Innovation Team (Grant No. T2023009), the Innovation Fund of the Hubei Key Laboratory of Intelligent Robots, Wuhan Institute of Technology (Grant No. HBIRL202409), and the Scientific Research Fund of Wuhan Institute of Technology (Grant No. 24QD07).

### v3

@@ IV-B2 Results on CAMERA25
Table II reports the performance of INKL-Pose compared with state-of-the-art methods on the CAMERA25 dataset. As shown in the table, INKL-Pose achieves outstanding accuracy across nearly all evaluation metrics. Although its IoU50 metric reaches 94.4%, slightly lower than SpherePose [21] (achieving 94.8%), it significantly outperforms all existing approaches on the remaining metrics, especially under the more rigorous pose accuracy thresholds. Specifically, compared to prior-based methods, INKL-Pose maintains a clear advantage without relying on pre-defined shape priors. In addition, when compared with AG-Pose [20], the current state-of-the-art shape-prior-free method, INKL-Pose achieves notable gains of 1.5% and 1.3% under the strict 5°2cm and 5°5cm metrics, respectively. These results, along with the strong performance observed on the REAL275 dataset, provide compelling evidence of the robustness and effectiveness of our proposed local-to-global geometric feature aggregation framework. Furthermore, INKL-Pose exhibits strong model efficiency with 16.7M parameters, ranking among the most lightweight methods evaluated except for AG-Pose [20] and GPV-Pose [35]. It also achieves a fast inference speed of 36 FPS on an NVIDIA RTX 4090D GPU, highlighting its potential for real-time applications.
@@
Figure 5: Visualization results of INKL-Pose on the HouseCat6D dataset. Green bounding boxes denote ground truth, while red bounding boxes represent the predictions from our INKL-Pose. The predicted boxes show high consistency with the ground truth.
IV-B3 Results on HouseCat6D
To further validate the generalization capability of the proposed INKL-Pose, we conduct experiments on the more challenging real-world dataset HouseCat6D [31]. As reported in Table III, INKL-Pose achieves state-of-the-art performance across all evaluation metrics. Compared with the current state-of-the-art GCE-Pose [37], our method surpasses it under all evaluation metric, with gains of 2.7% on IoU50, 0.4% on 5°2cm, 1.7% on 5°5cm, 2.1% on 10°2cm, and 3.5% on 10°5cm. Moreover, Fig. 5 presents qualitative results of the predicted poses, further showcasing the strong generalization ability of INKL-Pose in complex real-world scenarios.
Figure 6: Comparison of predicted keypoints with the baseline method in NOCS space. The first row shows the results of AG-Pose, the second row shows the results of AG-Pose with the proposed keypoint losses, and the third row shows the results of our INKL-Pose. Keypoint colors indicate their NOCS errors, which are linearly mapped within the range $[0,0.2]$ from green (error = 0) to red (error = 0.2).
IV-B4 Visualization of Predicted Keypoints
Fig. 6 compares the keypoints predicted by our method with those from the baseline method AG-Pose [20] in the NOCS space. To highlight the differences, keypoints are color-coded based on their NOCS errors, linearly mapped from green (error = 0) to red (error = 0.2). As shown in the figure, AG-Pose produces keypoints with higher NOCS errors and uneven surface coverage. Incorporating our keypoint loss functions into AG-Pose leads to moderate improvements in both accuracy and spatial distribution. In contrast, our INKL-Pose achieves the lowest NOCS errors and generates a more uniform and comprehensive keypoint layout that better captures the object’s surface geometry. For example, in the third column showing a mug with a handle, AG-Pose fails to localize the handle, while its variant with our losses partially recovers the region but still exhibits large errors. In comparison, our method accurately delineates the handle contour, resulting in more precise pose estimation.
@@ IV-C Ablation Studies
In this section, we conduct comprehensive ablation studies to evaluate the contributions of key components proposed in INKL-Pose. All experiments are performed on the REAL275 [10] dataset.
@@ IV-C1 Effect of the Number of Keypoints
TABLE IV: Effect of the number of keypoints. Default configurations are colored in gray, with best results highlighted in bold.
@@
w/ ($\mathcal{L}_{sep}$+$\mathcal{L}_{surf}$) 63.2 68.3 80.9 88.0
w/ ($\mathcal{L}_{div}$+$\mathcal{L}_{ocd}$) 60.9 67.0 79.9 87.2
w/ $\mathcal{L}_{sep}$ (Ours) 60.8 67.7 77.8 86.2
w/ $\mathcal{L}_{div}$ (AG-Pose) 59.1 66.9 77.2 86.0
w/ $\mathcal{L}_{surf}$ (Ours) 46.9 53.0 69.1 78.8
w/ $\mathcal{L}_{ocd}$ (AG-Pose) 11.8 18.3 24.7 45.6
@@
TABLE V: Effect of keypoint supervision. Default configurations are colored in gray, with best results highlighted in bold.
IV-C2 Effect of Keypoint Supervision
In the Instance-Adaptive Keypoint Detector (IAKD) module (Section III-C), we introduce two auxiliary loss terms, $\mathcal{L}_{surf}$ and $\mathcal{L}_{sep}$, which encourage uniform surface coverage and spatial diversity among keypoints. Table V summarizes the performance under different supervision strategies. As the table shows, without any keypoint supervision (“None”), the model performs poorly across all metrics, highlighting the necessity of proper guidance. When incorporating the surface coverage loss $\mathcal{L}_{ocd}$ proposed in AG-Pose, the performance further degrades, indicating that this loss is not well suited to our framework. In contrast, the proposed surface loss $\mathcal{L}_{surf}$ significantly improves performance over the unsupervised baseline.
For keypoint separation, both the $\mathcal{L}_{div}$ loss from AG-Pose and our proposed $\mathcal{L}_{sep}$ result in clear performance gains, although $\mathcal{L}_{sep}$ consistently achieves higher accuracy. Notably, the combination of $\mathcal{L}_{surf}$ and $\mathcal{L}_{sep}$ delivers the best performance across all metrics, demonstrating their complementary contributions to keypoint learning. These results confirm that the proposed losses more effectively guide the network to generate semantically consistent, geometrically meaningful, and spatially well-distributed keypoints, which ultimately enhances pose estimation performance.
IV-C3 Effect of Positional Embeddings
We further analyze the contribution of positional information to feature representation. In INKL-Pose, positional encodings are introduced at two distinct stages: the embedding $\mathcal{E}_{\mathcal{P}}$ in the Pointwise Feature Encoder (Section III-B), which compensates for the loss of absolute position cues, and $\mathcal{E}_{kpt}$ in the LKFA module (Section III-D), which enhances local spatial awareness. As reported in Table VI, removing either embedding individually leads to a noticeable performance drop across all metrics, highlighting their complementary roles in capturing geometric information.
@@ IV-C3 Effect of Positional Embeddings
TABLE VI: Effect of positional embeddings. Default configurations are colored in gray, with best results highlighted in bold.
IV-C4 Effect of the LKFA and GKFA Modules
We evaluate the contributions of the LKFA and GKFA modules introduced in Section III-D and Section III-E, respectively. As shown in Table VII, removing the LKFA results in a 2.1% drop in the stringent 5°2cm metric, primarily due to the loss of fine-grained geometric features. Furthermore, removing the GKFA leads to a larger decline of 3.0%, indicating its role in preserving global consistency. These results underscore that both local and global geometric information are crucial for establishing accurate correspondences on unseen objects.
@@ IV-C4 Effect of the LKFA and GKFA Modules
TABLE VII: Effect of the LKFA and GKFA modules. Default configurations are colored in gray, with best results highlighted in bold.
IV-C5 Effect of the Number of LKFA and GKFA Modules
We investigate the impact of stacking different numbers of keypoint feature aggregation blocks, each containing both a Local Keypoint Feature Aggregator (LKFA) and a Global Keypoint Feature Aggregator (GKFA). As shown in Table VIII, increasing the number of blocks from $S=4$ to $S=12$ consistently improves performance across all metrics, as deeper cascaded aggregation enhances both local and global keypoint representations. However, using $S=16$ blocks leads to performance degradation due to feature redundancy and oversmoothing, which reduce keypoint distinctiveness and impair pose estimation. Overall, $S=12$ achieves the best balance between accuracy and efficiency.
@@ IV-C5 Effect of the Number of LKFA and GKFA Modules
TABLE VIII: Effect of the number of LKFA and GKFA modules. Default configurations are colored in gray, with best results highlighted in bold.
IV-C6 Effect of Neighborhood Size
To evaluate the influence of neighborhood size $K$ in the LKFA module, we vary the number of nearest neighbors used during local feature aggregation. As shown in Table IX, when $K=1$, the model only uses the center point feature, which limits its ability to capture surrounding geometric and semantic context. Increasing $K$ enhances local structure modeling, with optimal performance achieved at $K=4$. Further increasing $K$ introduces redundant information, leading to no additional gains.
@@ IV-C6 Effect of Neighborhood Size
TABLE IX: Effect of neighborhood size. Default configurations are colored in gray, with best results highlighted in bold.
IV-C7 Effect of Bidirectional Mamba and Feature Sequence Flipping
We also examine the contributions of the bidirectional Mamba architecture and the proposed Feature Sequence Flipping (FSF) strategy within the GKFA module (Section III-E), as summarized in Table X. The results show that replacing the bidirectional Mamba with its unidirectional counterpart (Uni-Mamba) consistently degrades performance, confirming the advantage of bidirectional modeling for global feature aggregation. Additionally, we compare FSF with the commonly used Point Sequence Flipping (PSF) strategy. PSF exacerbates the non-causal nature of point clouds by reversing the sequence along spatial indices, disrupting the alignment between feature order and geometric structure, and leading to performance drops even below the unidirectional baseline. In contrast, FSF preserves the original point order while achieving bidirectional information flow in feature space, effectively improving robustness to the unordered and non-causal characteristics of point clouds.
@@ IV-C7 Effect of Bidirectional Mamba and Feature Sequence Flipping
TABLE X: Effect of bidirectional Mamba and feature sequence flipping. Default configurations are colored in gray, with best results highlighted in bold.
@@ V Conclusion
In this paper, we propose INKL-Pose, a novel keypoint-based framework for category-level object pose estimation that enhances generalization to unseen object instances through instance-adaptive keypoint learning and local-to-global geometric aggregation. Specifically, we introduce an Instance-Adaptive Keypoint Detector that predicts semantically consistent and geometrically informative keypoints, along with two complementary modules for feature refinement: a Local Keypoint Feature Aggregator that captures fine-grained geometric details and a Global Keypoint Feature Aggregator based on bidirectional Mamba with a Feature Sequence Flipping strategy for structure-aware global perception. Furthermore, we design a surface loss and a separation loss to explicitly supervise keypoint distribution, providing strong geometric guidance for robust pose estimation. Extensive experiments on CAMERA25, REAL275, and HouseCat6D demonstrate that INKL-Pose achieves state-of-the-art performance across both synthetic and real-world datasets, significantly outperforming existing approaches. In future work, we will focus on extending our framework to dynamic and temporal settings, enabling sequential pose tracking, and explore uncertainty-aware keypoint modeling to further improve robustness under occlusion and sensor noise.